### v2

@@ 6.2 Dominant Strategies
This paper focuses on ex-post strategies. However, all our results can be extended to dominant strategies as well. To extend Theorem 3.4 from an ex-post strategy not to shill to a dominant strategy is straight-forward: shill bidding in the Dutch auction always leads to 00 revenue, which means it is a weakly dominated strategy, regardless of what other bidders do. Further, since there exist no other auctions besides the Dutch auction that have an ex-post strategy not to shill bid, there can exist no other auction with a dominant strategy not to shill bid.
To extend Theorem 4.5 from an ex-post equilibrium to a dominant strategy equilibrium for real bidders, some care must be taken in considering the information sets of different bidders when they take actions. However, we can provide dominant-strategy equilibria versions of the English and ascending, screening auctions by assuming bidders move simultaneously each round of the English auction, as well in the second-price auction phase of the ascending, screening auction. Theorem 5.2 holds if we were to instead consider dominant strategies for real bidders as dominant strategy incentive compatibility is a stronger condition than ex-post incentive compatibility.
@@ Appendix A Model (Section 2) Appendix
$\Sigma_{S}=\left\{\sigma_{S}^{\prime}:\forall v_{-S},\exists v_{S}\text{ such that }\left(\sigma_{S}^{\prime},\sigma_{-S}(v_{-S})\right)=\sigma(v_{S},v_{-S};R=B)\right\}.$
@@
$\mathds{E}_{v_{-i}^{\prime},\tilde{R}}\left[u_{i}\left(\sigma\left(v_{i},v_{-i}^{\prime};\tilde{R}\right)\right)\right]\geq\mathds{E}_{v_{-i}^{\prime},\tilde{R}}\left[u_{i}\left(\sigma_{i}^{\prime}\left(v_{i},v_{-i}^{\prime};\tilde{R}\right),\sigma_{-i}\left(v_{i},v_{-i}^{\prime};\tilde{R}\right)\right)\right],$
and for all $\sigma_{S}^{\prime}\in\Sigma_{S}$,
$\mathds{E}_{v_{-S}^{\prime}}\left[u_{i}\left(\sigma\left(0,v_{-S}^{\prime};R\right)\right)\right]\geq\mathds{E}_{v_{-S}^{\prime}}\left[u_{i}\left(\sigma_{S}^{\prime}\left(0,v_{-S}^{\prime};R\right),\sigma_{-S}\left(0,v_{-S}^{\prime};R\right)\right)\right].$
In Definition A.1, we are restricting shill bidders to acting “as-if” they are real bidders by restricting their actions to those of real bidders with some valuation profile. This restriction allows us to move to a direct mechanism where shill-proofness is defined as it being an equilibrium (ex-interim for weak shill-proofness, ex-post for strong shill-proofness) for all shill bidders to report 00. Note that if we were to enlarge the set $\Sigma_{S}$ to be the set $\hat{\Sigma}_{S}$ of all strategy profiles, are main results would not change. For Theorems 3.4, 3.6 and 4.5, we are focused on augmented direct games and in such games $\Sigma_{S}=\hat{\Sigma}_{S}$. For Theorem 5.2, we know that $\Sigma_{S}\subset\hat{\Sigma}_{S}$ and our impossibility result must still hold if the set of possible deviations by shill bidders is larger; thus, the theorem still holds.
Lemma A.2.
An optimal auction $(G,\sigma)$ is winner-paying: For all $i$ and $v$,
$x_{i}\left(\sigma(v;B)\right)=0\implies t_{i}\left(\sigma(v;B)\right)=0.$
Proof.
By the ex-post IR constraint, when $x_{i}\left(\sigma(v;B)\right)=0$, we have $t_{i}\left(\sigma(v;B)\right)\leq 0.$ It then follows from the optimality that $t_{i}\left(\sigma(v;B)\right)=0$. To see this, note that for bidder $j\neq i$, equilibrium constraints on bidder $j$ slacken when moving from $t_{i}<0$ to $t_{i}=0$ and so her play will remain the same. Meanwhile from bidder $i$, the transfer strictly increases moving from $t_{i}<0$ to $t_{i}=0$. ∎
@@
    There exists $v$ such that $h\preceq z(\sigma(v;B))$.
@@
    If $h\notin Z$, then for $i=P(h)$, there exists $v_{i}$,$v_{i}^{\prime}$, and $v_{-i}$ such that
@@
      $h\prec z(\sigma(v_{i}^{\prime},v_{-i};B))$, and
@@
      $(x,t)\left(\sigma(v;B)\right)\neq(x,t)\left(\sigma(v_{i}^{\prime},v_{-i};B)\right)$.
@@
If $\left(G,\sigma\right)$ is a game equilibrium, then there exists a game equilibrium $\left(G^{\prime},\sigma^{\prime}\right)$ that is pruned and for all $v$, $(x,t)\left(\sigma(v;B)\right)=(x^{\prime},t^{\prime})\left(\sigma^{\prime}(v;B)\right)$.
@@
    There exists a direct mechanism $(\tilde{x},\tilde{t})$: for all $v$, $\tilde{x}(v)=x(\sigma(v;B))$ and $\tilde{t}(v)=t(\sigma(v;B)).$
@@
    There exists a choice menu rule $\mu$ that is a function of the potential values $V=V_{1}\times\cdots\times V_{N}$ and bidder $\xi$. This rule has an output of $L\geq 2$ choices characterized as $\left\{\left(W_{\ell},\vec{\xi}_{\ell}\right)\right\}_{\ell\in\left\{1,\ldots,L\right\}}$ where:
@@
      $\left\{W_{\ell}\right\}_{\ell\in L}$ forms a partition of $V_{\xi}$, $\vec{\xi}_{\ell}\in\left(B\cup\left\{\emptyset\right\}\right)\backslash\left\{\xi\right\}$, and $\vec{\xi}=\emptyset$ signifies the game has ended.
@@
      For any $\ell$ such that $\vec{\xi}_{\ell}\neq\emptyset$, let $\hat{V}^{\ell}=\left(V_{1},\ldots,V_{\xi-1},W_{\ell},V_{\xi+1},\ldots,V_{N}\right)$. Then, for any such $\ell$, there exists $v_{\xi_{\ell}},v_{\xi_{\ell}}^{\prime}\in\hat{V}^{\ell}_{\xi_{\ell}}$, and $v_{-\xi_{\ell}}\in\hat{V}^{\ell}_{-\xi_{\ell}}$ such that $(\tilde{x},\tilde{t})\left(v_{\xi_{\ell}},v_{-\xi_{\ell}}\right)\neq(\tilde{x},\tilde{t})\left(v_{\xi_{\ell}}^{\prime},v_{-\xi_{\ell}}\right).$ If $v_{\xi}\in W_{\ell}$, then the next player in the game is $\vec{\xi}_{\ell}$ and the menu presented to her is $\mu(\hat{V}^{\ell},\vec{\xi}_{\ell})$.
@@
      If $\ell$ is such that $\vec{\xi}_{\ell}=\emptyset$, then for all $v,v^{\prime}\in\hat{V}_{\ell}$, $\left(\tilde{x},\tilde{t}\right)\left(v\right)=\left(\tilde{x},\tilde{t}\right)\left(v^{\prime}\right).$
@@
      The first player to take an action is $\xi_{0}$, who is presented the menu $\mu(\mathcal{V}^{N},\xi_{0})$.
@@
To prove Condition i, we simply construct $(\tilde{x},\tilde{t})$ by iterating over all possible $v$ and defining $(\tilde{x},\tilde{t})$ as the outcome of $\sigma(v;B)$ in $G$.
To prove Condition ii, we first observe that by Definition A.1, shill bidders must act “as-if” they were real bidders and that we have restricted to pure strategies. Thus, we can always label actions as classes $\left(W_{\ell},\vec{\xi}_{\ell}\right)$ of a partition of the remaining possible values for the current player $\xi$ and satisfy Condition LABEL:*cond:rev_prin_2.LABEL:*cond:rev_prin_2_a. Condition LABEL:*cond:rev_prin_2.LABEL:*cond:rev_prin_2_c follows from the fact that $G$ is well defined (with each terminal history associated with a single outcome). Condition LABEL:*cond:rev_prin_2.LABEL:*cond:rev_prin_2_d is simply mapping the first player in $G$ to $\xi_{0}$ and the auctioneer has no information on bidders’ values yet. The fact that $L\geq 2$ is equivalent to Conditions i and ii of Definition A.3, and Condition iii of Definition A.3 is equivalent to Condition LABEL:*cond:rev_prin_2.LABEL:*cond:rev_prin_2_b here. We can then apply Lemma A.4 to find a game that satisfies these properties. ∎
Appendix B Dutch Auctions (Section 3) Appendix
B.1 Strongly Shill-Proof Auctions (Section 3.1) Appendix
@@ Appendix B Dutch Auctions (Section 3) Appendix
Towards contradiction, suppose there exists a strongly shill-proof auction $(\tilde{x}^{*},\tilde{t},\mu,\xi_{0})$, player $\xi$, and values $v_{\xi},v_{-\xi},v_{-\xi}^{\prime}$ such that $\tilde{x}_{\xi}(v)=\tilde{x}_{\xi}(v_{\xi},v_{\xi}^{\prime})$, but $\tilde{t}_{\xi}(v)\neq\tilde{t}_{\xi}(v_{\xi},v_{\xi}^{\prime})$. WLOG, suppose $\tilde{t}_{\xi}(v)>\tilde{t}_{\xi}(v_{\xi},v_{\xi}^{\prime})$. Now by Lemma A.2, $\xi$ can only have two different transfers if that player wins the item under the allocation. Then, take $R=\left\{\xi\right\}$ and by monotonicity, $\tilde{t}_{\xi}(v)>\tilde{t}_{\xi}(v_{\xi},v_{\xi}^{\prime})\geq\tilde{t}(v_{\xi}^{\prime};0)$ and thus shilling increases revenue and the auction is not strongly shill-proof. ∎
Given our auction $(\tilde{x},\tilde{t},\mu,\xi_{0})$, we define $X_{i}(v_{i};V)$ and $T_{i}(v_{i};V)$ to be the ex-interim quantity and transfer rules, respectively, when bidder $i$ has value $v_{i}$ and the set of potential values for all bidders is $V=V_{1}\times\cdots\times V_{N}$.
@@
For every optimal, weakly shill-proof auction $(\tilde{x}^{*},\tilde{t},\mu,\xi_{0})$, the ex-interim transfer rule for bidder $i$ is
$T_{i}(v_{i};V)=X_{i}(v_{i};V)v_{i}-\sum_{m:v^{j_{m}}<v_{i}}\left[X_{i}(v^{j_{m}};V)\cdot(v^{j_{m+1}}-v^{j_{m}})\right],$
where $\left\{v^{j_{m}}\right\}_{m}$ are the ordered atoms of $V_{i}$.
@@
To prove that $T$ has the claimed form, we will consider a specific non-truthful reporting: if a bidder has value $v^{m}$, she commits to mis-reporting (selecting partitions) $v^{m^{\prime}}$ for the rest of the game. We now follow the proof of Theorem 1 of Elkind (2007). Since our direct mechanism is an equilibrium for real bidders, we must have
$X_{i}(v^{j_{m}};V)v^{j_{m}}-T_{i}(v^{j_{m}};V)\geq X_{i}(v^{j_{m-1}};V)v^{j_{m}}-T_{i}(v^{j_{m-1}};V),\text{ and}$
$X_{i}(v^{j_{m-1}};V)v^{j_{m-1}}-T_{i}(v^{j_{m-1}};V)\geq X_{i}(v^{j_{m}};V)v^{j_{m-1}}-T_{i}(v^{j_{m}};V).$
@@
$U_{i}(v^{j_{m}};V)\geq U_{i}(v^{j_{m-1}};V)+(v^{j_{m}}-v^{j_{m-1}})X_{i}(v^{j_{m-1}};V),\text{ and}$
$U_{i}(v^{j_{m-1}};V)\geq U_{i}(v^{j_{m}};V)-(v^{j_{m}}-v^{j_{m-1}})X_{i}(v^{j_{m}};V).$
Thus, $(v^{j_{m}}-v^{j_{m-1}})X_{i}(v^{j_{m-1}};V)\leq U_{i}(v^{j_{m}};V)-U_{i}(v^{j_{m-1}};V)\leq(v^{j_{m}}-v^{j_{m-1}})X_{i}(v^{j_{m}};V).$ Hence, any IC mechanism is such that
$U_{i}(v^{j_{m}};V)$ $=U_{i}(v^{j_{1}};V)+\sum_{k=2}^{m}(v^{j_{m}}-v^{j_{m-1}})\tilde{X}_{i}(v^{j_{m}};V)$
where $\text{where }\tilde{X}_{i}(v^{j_{m}};V)\in\left[X_{i}(v^{j_{m-1}};V),X_{i}(v^{j_{m}};V)\right].$
@@
$T_{i}(v^{j_{m}};V)=X_{i}(v^{j_{m}};V)v^{j_{m}}-U_{i}(v^{j_{1}};V)-\sum_{k=2}^{m}(v^{j_{m}}-v^{j_{m-1}})\tilde{X}_{i}(v^{j_{m}};V).$ (2)
By the ex-post IR condition, we have $U_{i}(v^{j_{1}};V)\geq 0$ for all $V$. So, solving for the optimal transfer rule from Equation 2,
$T_{i}^{*}(v^{j_{m}};V)$ $=\max_{U_{i},\tilde{X}}\left[X_{i}(v^{j_{m}};V)v^{j_{m}}-U_{i}(v^{j_{1}};V)-\sum_{k=2}^{m}(v^{j_{m}}-v^{j_{m-1}})\tilde{X}_{i}(v^{j_{m}};V)\right]$
[…]
The solution to this maximization is $U_{i}(v^{j_{1}};V)=0,\tilde{X}_{i}(v^{j_{m}};V)=X_{i}(v^{j_{m-1}};V).$ Thus, Equation 2 becomes
$T_{i}(v_{i};V)$ $=X_{i}(v_{i};V)v_{i}-\sum_{m:v^{j_{m}}<v_{i}}\left[X_{i}(v^{j_{m}};V)\cdot(v^{j_{m+1}}-v^{j_{m}})\right].$
@@
Consider a strongly shill-proof, public, optimal auction $(\tilde{x}^{*},\tilde{t},\mu,\xi_{0})$. Fix $V,\xi$ and consider any $(W,\vec{\xi})\in\mu(V,\xi)$. If there exists $v,v^{\prime}\in V$ such that $v_{\xi},v_{\xi}^{\prime}\in W$ and $\mu(V,\xi)$ is the last action $\xi$ takes, then,
$\tilde{x}^{*}_{\xi}(v)=\tilde{x}^{*}_{\xi}(v^{\prime})=1\implies\tilde{t}_{\xi}(v)=\tilde{t}_{\xi}(v^{\prime})=\frac{T_{\xi}(\underline{W};V)}{X_{\xi}(\underline{W};V)},$
@@
Since an auction cannot distinguish between values in the same choice set, we can apply Lemma 3.2 to conclude that if $\mu(V,\xi)$ is the last action $\xi$ takes, then $\tilde{t}(v)=\tilde{t}(v^{\prime})$. To conclude the proof, we note that $\xi$ wins no matter what her value is in $W$ and then apply Lemma A.2 to observe that $T_{\xi}(\underline{W};V)=\tilde{t}_{\xi}(v)\cdot X_{\xi}(\underline{W};V)$. ∎
@@
We first show that the Dutch auction is a well-defined, i.e., that the stopping rule allows for the auction to be orderly and optimal. [^34] The auction is public by definition. We then show that it is strongly shill-proof. Finally, we show that there are no other public, strongly shill-proof, orderly, optimal auctions.
@@
The Dutch Auction is Orderly, Optimal, and Strongly Shill-Proof. The Dutch auction quantity and transfer rule are orderly and optimal (as well as ex-post IR and monotone). Indeed, by construction, the next player $\vec{\xi}$ is always the player with the potentially highest value (including for tie-breaking). So, if that player indicates that she is of the highest possible type, the outcome (allocation and transfer) is fully determined and the auction ends. The auction ends once there are no players who could have values weakly greater than $\rho^{*}$.

### v3

@@ 6.2 Dominant Strategies
This paper focuses on ex-post strategies. However, all our results can be extended to dominant strategies as well. To extend Theorem 3.6 from an ex-post strategy not to shill to a dominant strategy is straight-forward: shill bidding in the Dutch auction always leads to 0 revenue, which means it is a weakly dominated strategy, regardless of what other bidders do. Further, since there exist no other auctions besides the Dutch auction that have an ex-post strategy not to shill bid, there can exist no other auction with a dominant strategy not to shill bid.
To extend Proposition 4.12 from an ex-post equilibrium to a dominant strategy equilibrium for real bidders, some care must be taken in considering the information sets of different bidders when they take actions. However, we can provide dominant-strategy equilibria versions of the English and ascending, screening auctions by assuming bidders move simultaneously each round of the English auction, as well in the second-price auction phase of the ascending, screening auction. Theorem 4.3 holds if we were to instead consider dominant strategies for real bidders as dominant strategy incentive compatibility is a stronger condition than ex-post incentive compatibility.
@@ Appendix A Model (Section 2) Appendix
$\Sigma_{S}=\left\{\sigma_{S}^{\prime}:\forall\theta_{-S},\exists\theta_{S}\text{ such that }\left(\sigma_{S}^{\prime},\sigma_{-S}(\theta_{-S})\right)=\sigma(\theta_{S},\theta_{-S};R=B)\right\}.$
@@
[…]
and for all $i\in S$ and $\sigma_{S}^{\prime}\in\Sigma_{S}$,
$\mathds{E}_{\theta_{-S}^{\prime}}\left[u_{i}\left(\sigma\left(0,\theta_{-S}^{\prime};R\right)\right)\right]\geq\mathds{E}_{\theta_{-S}^{\prime}}\left[u_{i}\left(\sigma_{S}^{\prime}\left(0,\theta_{-S}^{\prime};R\right),\sigma_{-S}\left(0,\theta_{-S}^{\prime};R\right)\right)\right].$
In Definition A.1, the strategy space $\Sigma_{S}$ is restricting shill bidders to only take actions that could be played on-path by a real bidder. In other words, shill bidders must act “as-if” they are real bidders and not take any actions that would definitively prove that they are shill bidders. This restriction allows us to move to a direct mechanism where shill-proofness is defined as it being an equilibrium (ex-interim for weak shill-proofness, ex-post for strong shill-proofness) for all shill bidders to report 0. Note that if we were to enlarge the strategy space $\Sigma_{S}$ to be the set $\hat{\Sigma}_{S}$ of all strategy profiles, are main results would not change. For Theorem 3.6, Proposition 4.6, and Proposition 4.12, we are focused on augmented direct games and in such games $\Sigma_{S}=\hat{\Sigma}_{S}$. For Theorem 4.3, we know that $\Sigma_{S}\subset\hat{\Sigma}_{S}$ and our impossibility result must still hold if the set of possible deviations by shill bidders is larger; thus, the theorem still holds.
@@
    There exists $\theta$ such that $h\preceq z(\sigma(\theta;B))$.
@@
    If $h\notin Z$, then for $i=P(h)$, there exists $\theta_{i}$$\theta_{i}^{\prime}$, and $\theta_{-i}$ such that
@@
      $h\prec z(\sigma(\theta_{i}^{\prime},\theta_{-i};B))$, and
@@
      $(x,t)\left(\sigma(\theta;B)\right)\neq(x,t)\left(\sigma(\theta_{i}^{\prime},\theta_{-i};B)\right)$.
@@
If $\left(G,\sigma\right)$ is a game equilibrium, then there exists a game equilibrium $\left(G^{\prime},\sigma^{\prime}\right)$ that is pruned and for all $\theta$, $(x,t)\left(\sigma(\theta;B)\right)=(x^{\prime},t^{\prime})\left(\sigma^{\prime}(\theta;B)\right)$.
@@
[…]
@@
    There exists a choice menu rule $\mu$ that is a function of the potential values $\Theta=\Theta_{1}\times\cdots\times\Theta_{N}$ and bidder $\xi$. This rule has an output of $L\geq 2$ choices characterized as $\left\{\left(W_{\ell},\vec{\xi}_{\ell}\right)\right\}_{\ell\in\left\{1,\ldots,L\right\}}$ where:
@@
[…]
@@
      For any $\ell$ such that $\vec{\xi}_{\ell}\neq\emptyset$, let $\hat{\Theta}^{\ell}=\left(\Theta_{1},\ldots,\Theta_{\xi-1},W_{\ell},\Theta_{\xi+1},\ldots,\Theta_{N}\right)$. Then, for any such $\ell$, there exists $\theta_{\xi_{\ell}},\theta_{\xi_{\ell}}^{\prime}\in\hat{\Theta}^{\ell}_{\xi_{\ell}}$, and $\theta_{-\xi_{\ell}}\in\hat{\Theta}^{\ell}_{-\xi_{\ell}}$ such that $(\tilde{x},\tilde{t})\left(\theta_{\xi_{\ell}},\theta_{-\xi_{\ell}}\right)\neq(\tilde{x},\tilde{t})\left(\theta_{\xi_{\ell}}^{\prime},\theta_{-\xi_{\ell}}\right).$ If $\theta_{\xi}\in W_{\ell}$, then the next player in the game is $\vec{\xi}_{\ell}$ and the menu presented to her is $\mu(\hat{\Theta}^{\ell},\vec{\xi}_{\ell})$.
@@
      If $\ell$ is such that $\vec{\xi}_{\ell}=\emptyset$, then for all $\theta,\theta^{\prime}\in\hat{\Theta}_{\ell}$, $\left(\tilde{x},\tilde{t}\right)\left(\theta\right)=\left(\tilde{x},\tilde{t}\right)\left(\theta^{\prime}\right).$
@@
      The first player to take an action is $\xi_{0}$, who is presented the menu $\mu(\vartheta^{N},\xi_{0})$.
@@
[…]
To prove Condition ii, we first observe that by Definition A.1, shill bidders must act “as-if” they were real bidders and that we have restricted to pure strategies. Thus, we can always label actions as classes $\left(W_{\ell},\vec{\xi}_{\ell}\right)$ of a partition of the remaining possible values for the current player $\xi$ and satisfy Condition iia. The fact that $L\geq 2$ is equivalent to Conditions i and ii of Definition A.2, and Condition iii of Definition A.2 is equivalent to Condition iib here. We can then apply Lemma A.3 to find a game that satisfies these properties. Condition iic follows from the fact that $G$ is well-defined (with each terminal history associated with a single outcome). Condition iid is simply mapping the first player in $G$ to $\xi_{0}$ and the auctioneer has no information on bidders’ values yet. ∎
Appendix B Strongly Shill-Proof Auctions (Section 3) Appendix
Lemma B.1.
An optimal auction $(G,\sigma)$ is winner-paying: For all $i$ and $v$,
$x_{i}\left(\sigma(\theta;B)\right)=0\implies t_{i}\left(\sigma(\theta;B)\right)=0.$
Proof.
By the ex-post IR constraint, when $x_{i}\left(\sigma(\theta;B)\right)=0$, we have $t_{i}\left(\sigma(\theta;B)\right)\leq 0.$ It then follows from optimality that $t_{i}\left(\sigma(\theta;B)\right)=0$. To see this, note that for bidder $j\neq i$, equilibrium constraints on bidder $j$ slacken when moving from $t_{i}<0$ to $t_{i}=0$ and so her play will remain the same. Meanwhile the transfer from bidder $i$ strictly increases moving from $t_{i}<0$ to $t_{i}=0$. ∎
@@ Appendix B Strongly Shill-Proof Auctions (Section 3) Appendix
Towards contradiction, suppose there exists a strongly shill-proof auction $(\tilde{x},\tilde{t},\mu,\xi_{0})$, player $\xi$, and values $\theta_{\xi},\theta_{-\xi},\theta_{-\xi}^{\prime}$ such that $\tilde{x}_{\xi}(\theta)=\tilde{x}_{\xi}(\theta_{\xi},\theta_{-\xi}^{\prime})$, but $\tilde{t}_{\xi}(\theta)\neq\tilde{t}_{\xi}(\theta_{\xi},\theta_{-\xi}^{\prime})$. WLOG, suppose $\tilde{t}_{\xi}(\theta)>\tilde{t}_{\xi}(\theta_{\xi},\theta_{-\xi}^{\prime})$. Now by Lemma B.1, $\xi$ can only have two different transfers if that player wins the item under the allocation. Then, take $R=\left\{\xi\right\}$ and by monotonicity, $\tilde{t}_{\xi}(\theta)>\tilde{t}_{\xi}(\theta_{\xi},\theta_{-\xi}^{\prime})\geq\tilde{t}(\theta_{\xi},0)$ and thus shilling increases revenue and the auction is not strongly shill-proof. ∎
Given our auction $(\tilde{x},\tilde{t},\mu,\xi_{0})$, we define $X_{i}(\theta_{i};\Theta)$ and $T_{i}(\theta_{i};\Theta)$ to be the ex-interim quantity and transfer rules, respectively, when bidder $i$ has value $\theta_{i}$ and the set of potential values for all bidders is $\Theta$.
@@
For every optimal auction $(\tilde{x}^{*},\tilde{t},\mu,\xi_{0})$, the ex-interim transfer rule for bidder $i$ is
$T_{i}(\theta_{i};\Theta)=X_{i}(\theta_{i};\Theta)\theta_{i}-\sum_{m:\theta^{j_{m}}<\theta_{i}}\left[X_{i}(\theta^{j_{m}};\Theta)\cdot(\theta^{j_{m+1}}-\theta^{j_{m}})\right],$
where $\left\{\theta^{j_{m}}\right\}_{m}$ are the ordered atoms of $\Theta_{i}$.
@@
To prove that $T$ has the claimed form, we will consider a specific non-truthful reporting: if a bidder has value $\theta^{m}$, she commits to mis-reporting (selecting partitions) $\theta^{m^{\prime}}$ for the rest of the game. We now follow the proof of Theorem 1 of Elkind (2007). Since our direct mechanism is an equilibrium for real bidders, we must have that
$X_{i}(\theta^{j_{m}};\Theta)\theta^{j_{m}}-T_{i}(\theta^{j_{m}};\Theta)\geq X_{i}(\theta^{j_{m-1}};\Theta)\theta^{j_{m}}-T_{i}(\theta^{j_{m-1}};\Theta),\text{ and}$
$X_{i}(\theta^{j_{m-1}};\Theta)\theta^{j_{m-1}}-T_{i}(\theta^{j_{m-1}};\Theta)\geq X_{i}(\theta^{j_{m}};\Theta)\theta^{j_{m-1}}-T_{i}(\theta^{j_{m}};\Theta).$
@@
$U_{i}(\theta^{j_{m}};\Theta)\geq U_{i}(\theta^{j_{m-1}};\Theta)+(\theta^{j_{m}}-\theta^{j_{m-1}})X_{i}(\theta^{j_{m-1}};\Theta),\text{ and}$
$U_{i}(\theta^{j_{m-1}};\Theta)\geq U_{i}(\theta^{j_{m}};\Theta)-(\theta^{j_{m}}-\theta^{j_{m-1}})X_{i}(\theta^{j_{m}};\Theta).$
Thus, $(\theta^{j_{m}}-\theta^{j_{m-1}})X_{i}(\theta^{j_{m-1}};\Theta)\leq U_{i}(\theta^{j_{m}};\Theta)-U_{i}(\theta^{j_{m-1}};\Theta)\leq(\theta^{j_{m}}-\theta^{j_{m-1}})X_{i}(\theta^{j_{m}};\Theta).$ Hence, any IC mechanism is such that
$U_{i}(\theta^{j_{m}};\Theta)$ $=U_{i}(\theta^{j_{1}};\Theta)+\sum_{k=2}^{m}(\theta^{j_{m}}-\theta^{j_{m-1}})\tilde{X}_{i}(\theta^{j_{m}};\Theta)$
$\text{where }\tilde{X}_{i}(\theta^{j_{m}};\Theta)\in\left[X_{i}(\theta^{j_{m-1}};\Theta),X_{i}(\theta^{j_{m}};\Theta)\right].$
@@
$T_{i}(\theta^{j_{m}};\Theta)=X_{i}(\theta^{j_{m}};\Theta)\theta^{j_{m}}-U_{i}(\theta^{j_{1}};\Theta)-\sum_{k=2}^{m}(\theta^{j_{m}}-\theta^{j_{m-1}})\tilde{X}_{i}(\theta^{j_{m}};\Theta).$ (2)
By the ex-post IR condition, we have $U_{i}(\theta^{j_{1}};\Theta)\geq 0$ for all $\Theta$. So, solving for the optimal transfer rule from Equation 2,
$T_{i}^{*}(\theta^{j_{m}};\Theta)$ $=\max_{U_{i},\tilde{X}}\left[X_{i}(\theta^{j_{m}};\Theta)\theta^{j_{m}}-U_{i}(\theta^{j_{1}};\Theta)-\sum_{k=2}^{m}(\theta^{j_{m}}-\theta^{j_{m-1}})\tilde{X}_{i}(\theta^{j_{m}};\Theta)\right]$
$\text{ such that }U_{i}(\theta^{j_{1}};\Theta)\geq 0\text{ and }\tilde{X}_{i}(\theta^{j_{m}};\Theta)\in\left[X_{i}(\theta^{j_{m-1}};\Theta),X_{i}(\theta^{j_{m}};\Theta)\right].$
The solution to this maximization is $U_{i}(\theta^{j_{1}};\Theta)=0,\tilde{X}_{i}(\theta^{j_{m}};\Theta)=X_{i}(\theta^{j_{m-1}};\Theta).$ Thus, Equation 2 becomes
$T_{i}(\theta_{i};\Theta)$ $=X_{i}(\theta_{i};\Theta)\theta_{i}-\sum_{m:\theta^{j_{m}}<\theta_{i}}\left[X_{i}(\theta^{j_{m}};\Theta)\cdot(\theta^{j_{m+1}}-\theta^{j_{m}})\right].$
@@
Consider a strongly shill-proof and optimal auction $(\tilde{x}^{*},\tilde{t},\mu,\xi_{0})$. Fix $\Theta,\xi$ and consider any $(W,\vec{\xi})\in\mu(\Theta,\xi)$. If there exists $\theta,\theta^{\prime}\in\Theta$ such that $\theta_{\xi},\theta_{\xi}^{\prime}\in W$ and $\mu(\Theta,\xi)$ is the last action $\xi$ takes, then,
$\tilde{x}^{*}_{\xi}(\theta)=\tilde{x}^{*}_{\xi}(\theta^{\prime})=1\implies\tilde{t}_{\xi}(\theta)=\tilde{t}_{\xi}(\theta^{\prime})=\frac{T_{\xi}(\underline{W};\Theta)}{X_{\xi}(\underline{W};\Theta)},$
@@
Since an auction cannot distinguish between values in the same choice set, we can apply Lemma 3.1 to conclude that if $\mu(\Theta,\xi)$ is the last action $\xi$ takes, then $\tilde{t}(\theta)=\tilde{t}(\theta^{\prime})$. To conclude the proof, we note that $\xi$ wins no matter what her value is in $W$ and then apply Lemma B.1 to conclude that $T_{\xi}(\underline{W};\Theta)=\tilde{t}_{\xi}(\theta)\cdot X_{\xi}(\underline{W};\Theta)$. ∎
Proof of Proposition 3.3
Under a static information structure, the single-action, first-price auction directly corresponds to the direct mechanism. In a pay-as-bid mechanism, a shill bidder’s actions have no way to influence the payments except by changing the allocation and by the definition of orderly, if a shill bidder changes the allocation she wins the auction, which is dominated. To prove that any more informative structure is not strongly shill-proof, consider some $i,j$ and $\theta$ where $\mathcal{I}(h)\neq\mathcal{I}(h_{0})$. By definition, information sets form a partition and so $\mathcal{I}(h)\cap\mathcal{I}(h_{0})=\emptyset$. Observe that $j$ can have any type at both $h$ and $h_{0}$ given that each bidder takes a single action. Now, consider the minimal $m$ such that $j$ believes ex-interim that she could win with type $\theta^{m}$. By assumption, $\theta^{m}>\theta^{\rho^{*}}$. Because we are considering strong shill-proofness, we can assume without loss that $j$ does in fact win the auction at $\theta^{m}$. So, at $h$, we can then apply Lemma B.3 to complete the proof that $j$ must pay more at $h$ than at $h_{0}$ and so if $i\in S$, shill bidding is profitable at this type realization. ∎
@@
We first show that the Dutch auction with reserve price $\theta^{\rho}$ is a well-defined, i.e., that the stopping rule allows for the auction to be orderly. [^31] The auction is public by definition. We then show that it is strongly shill-proof. Finally, we show that there are no other public, strongly shill-proof, orderly, optimal auctions.
@@
The Dutch Auction is Orderly, Optimal, and Strongly Shill-Proof. The Dutch auction quantity rule is orderly and the transfer rule is ex-post IR and monotone. Indeed, by construction, the next player $\vec{\xi}$ is always the player with the potentially highest value (including for tie-breaking). So, if that player indicates that she is of the highest possible type, the outcome (allocation and transfer) is fully determined and the auction ends. The auction ends once there are no players who could have values weakly greater than $\theta^{\rho}$.